### v1

# Sample Path Regularity of Gaussian Processes from the Covariance Kernel

Abstract
Gaussian processes (GPs) are the most common formalism for defining probability distributions over spaces of functions. While applications of GPs are myriad, a comprehensive understanding of GP sample paths, i.e. the function spaces over which they define a probability measure on, is lacking. In practice, GPs are not constructed through a probability measure, but instead through a mean function and a covariance kernel. In this paper we provide necessary and sufficient conditions on the covariance kernel for the sample paths of the corresponding GP to attain a given regularity. We use the framework of Hölder regularity as it grants us particularly straightforward conditions, which simplify further in the cases of stationary and isotropic GPs. We then demonstrate that our results allow for novel and unusually tight characterisations of the sample path regularities of the GPs commonly used in machine learning applications, such as the Matérn GPs.
Keywords: Gaussian processes, Gaussian random fields, differentiability, Hölder regularity, Sobolev regularity, stationary kernels, isotropic kernels, Matérn kernels
1 Introduction
Gaussian processes (GPs) provide a formalism to assign probability distributions over spaces of functions. That distribution is principally characterised and controlled by the process’ covariance function, which is a positive definite kernel. Such kernels are also associated with reproducing kernel Hilbert spaces (RKHSs). However, it is relatively widely known that the sample paths of the associated GP are not generally elements of the RKHS, but form a “larger” space of typically less regular functions (Kanagawa et al., 2018, Section 4). This sample path space is much harder to characterise than the RKHS.
GP regression is widely applied in statistics and machine learning for inference from physical observations. In such settings, practitioners mostly concern themselves with the posterior mean function (which is an element of the RKHS) and the marginal variance (which is not in the RKHS, but inherits its regularity) while largely ignoring the regularity of the sample paths and other properties of the support of the prior probability measure. But recently relevant use cases for GPs in computational tasks urgently require a more careful, and ideally tight analysis of the sample path regularity, which we provide in this work. For instance, Pförtner et al. (2023)author=Marvin,color=yellowauthor=Marvin,color=yellowtodo: author=Marvin,color=yellowThis should probably be [Pförtner et al., 2022] instead of 2023 recently showed that a large class of classic solution methods for linear partial differential equations (PDEs) can be interpreted as computing the posterior mean of specific Gaussian processes, i.e. as base instances of probabilistic numerical methods. To infer the solution to a PDE using a probabilistic numerical method, we want to construct a Gaussian process prior for it. It should be tailored to the problem as tightly as possible, for the following reasons:
  1. 1.
    We need to ensure that the PDE’s differential operator is well-defined on all samples of the GP. Hence, the sample paths must be regular enough. Here, regularity typically refers to the existence of a number of strong or weak partial derivatives.
  2. 2.
    But we also want the GP posterior to be a useful uncertainty estimate over the solution of the PDE. Hence, we do not want to needlessly impose additional regularity constraints on the sample paths. The sample paths should be as irregular as possible, to avoid overconfidence. Such cautious models may also be advisable on numerical grounds, to avoid instabilities such as Gibbs or Runge phenomena.
The above is just one example of a setting in which one would like to characterise the regularity of GP sample paths as tightly as possible. Gaussian process models are increasingly common in machine learning applications. Further relevant examples beyond GP regression include generalised coordinates and generalised Bayesian filtering (Friston et al. (2010) and Heins et al. (2023)). These are generic methods for analysis, filtering and control of stochastic differential equations driven by signals that admit a high number of derivatives. For analytic and practical tractability, the noise therein is usually modelled with GPs. Characterisations of GP sample path regularity are therefore crucial for understanding in which situations these methods can be applied.
1.1 Summary of contributions
A simplified consequence of our main result, Theorem 7, may be written as follows:
Corollary 1
Let $k\colon\mathbb{R}^{d}\times\mathbb{R}^{d}\to\mathbb{R}$ be a symmetric positive definite kernel. If,
  • for general $k$, all partial derivatives of the form $\frac{\partial^{2n}k}{\partial\bm{x}^{\bm{\alpha}}\partial\bm{y}^{\bm{\alpha}}}$ for multi-indices $\bm{\alpha}\in\mathbb{N}_{0}^{d}$ with $|\bm{\alpha}|=n$ exist and are (locally) Lipschitz,
  • for stationary $k(\bm{x},\bm{y})=k_{\delta}(\bm{x}-\bm{y})$, $\frac{\partial^{2n}k_{\delta}}{\partial\bm{x}^{\alpha}}$ for $|\bm{\alpha}|=n$ exist and are (locally) Lipschitz,
  • for isotropic $k(\bm{x},\bm{y})=k_{r}(\|\bm{x}-\bm{y}\|)$, $k_{r}^{(2n)}$ exists and is (locally) Lipschitz,
then the sample paths of $f\sim\mathcal{GP}(0,k)$ are $n$ times continuously differentiable.
Note that, by the mean value theorem, the existence of one additional continuous derivative of $k$ is sufficient for local Lipschitz continuity of the lower derivatives.
Theorem 7 is sharper than this. It provides necessary and sufficient conditions on the regularity of the kernel for the sample paths to attain a given Hölder regularity.author=Marvin,color=yellowauthor=Marvin,color=yellowtodo: author=Marvin,color=yellowMention that Hölder regularity is in itself important for e.g. strong PDE applications? Just as the above corollary, the theorem encompasses statements for stationary and isotropic GPs, which are of great importance in applications.
We apply Theorem 7 in particular to the Matérn and Wendland GPs to obtain Propositions 9 and 11 respectively, consequences of which are the following:
Corollary 2
The sample paths of a centered Matérn GP with smoothness parameter $\nu\not\in\mathbb{N}_{0}$ are exactly $\lfloor\nu\rfloor$ times continuously differentiable. In particular, when $\nu=n+1/2$ for some $n\in\mathbb{N}_{0}$, the sample paths are exactly $n$ times continuously differentiable.
Corollary 3
The sample paths of a centered Wendland GP with degree parameter $n$ are exactly $n$ times continuously differentiable.
We then describe in Proposition 13 how to apply Theorem 7 to GPs with tensor product covariance kernels.
In Theorem 16 we summarise known results about general and stationary GP sample path Sobolev regularity, specialise them to isotropic GPs and compare them to Theorem 7. We see that in practice one should not expect to obtain a greater number of weak (Sobolev) derivatives than of strong derivatives for GP sample paths. This provides further evidence that Theorem 7 contains all that is needed in practice.
Let us finally note that all regularity properties considered in this paper are local, and we do not aim at capturing global properties of sample paths in this work.
1.2 Related work
Some of the earliest work on sample path regularity of GPs may be attributed to Fernique (1975). In that work, necessary and sufficient conditions for sample path continuity of stationary GPs are proved. Since then, many authors have extended these continuity results. Potthoff (2009) provides an approachable overview of how such results can be obtained. It also investigates conditions for uniform continuity, which is a global regularity property. Azmoodeh et al. (2014), which we rely on in this work, provides necessary and sufficient conditions for Hölder continuity of GPs. This can be seen as an application of the Kolmogorov continuity theorem to GPs, as well as its converse.
Differentiability has also been investigated by many authors, including Scheuerer (2010b) for general random fields, Adler and Taylor (2007), Potthoff (2010) and Henderson (2024) for Sobolev regularity. The sample path regularity of Matérn GPs was studied in Scheuerer (2010a) (see specifically Scheuerer (2010a, Examples 5.3.19, 5.5.6 & 5.5.12)) and of tensor product of one dimensional Matérn GPs in Wang et al. (2021).
Compared to the existing literature, the present work provides general, versatile and easily applicable results for the study of GP sample path continuity and differentiability, without sacrificing tightness.
2 Preliminaries
Definition 4
A Gaussian process on a set $O$ is a map $f\colon O\times\Omega\to\mathbb{R}$, where $\Omega$ is a probability space, such that for all $x\in O$, $f(x,\cdot)\colon O\to\mathbb{R}$ is measurable, and such that for each $X:=(x_{1},\dots,x_{N})\in O^{N}$, the map $f(X,\cdot)\colon\Omega\to\mathbb{R}^{N}$ given by $f(X,\omega)=(f(x_{1},\omega),\dots,f(x_{N},\omega))$ is a multivariate Gaussian random variable.
The maps $f(\cdot,\omega)\colon O\to\mathbb{R}$ for $\omega\in\Omega$ are the sample paths of the GP.
We would like to study the regularity of the sample paths of a GP. However, Definition 4 is not convenient to work with in practice: one rarely constructs the probability space $\Omega$ of a GP. Instead, one characterises a GP by its mean and its covariance kernel.
Definition 5
The mean $\mu$ of a GP $f\colon O\times\Omega\to\mathbb{R}$ is the map
$\mu\colon O\to\mathbb{R},x\mapsto\mathbb{E}(f(x,\cdot)).$
The covariance kernel $k$ is the map
$k\colon O\times O\to\mathbb{R},(x,y)\mapsto\operatorname{cov}(f(x,\cdot),f(y,\cdot)).$
The GP $f$ is said to be centered if $\mu=0$.
Conversely, given maps $\mu\colon O\to\mathbb{R}$ and $k\colon O\times O\to\mathbb{R}$ with $k$ symmetric positive definite – meaning that for any finite set of points $\{x_{1},\dots,x_{N}\}\subset\Omega$, the matrix $(k(x_{i},x_{j}))_{i,j}$ is symmetric positive semi-definite – one can construct a probability space $\Omega$, and a GP $f\colon O\times\Omega\to\mathbb{R}$ with mean $\mu$ and covariance kernel $k$ (Klenke, 2014, Theorem 14.36). We then write $f\sim\mathcal{GP}(\mu,k)$.
At this point it is important to note that $f\sim\mathcal{GP}(\mu,k)$ does not uniquely specify $f$ or the probability space $\Omega$. However it uniquely specifies the finite dimensional distributions of $f$ – since multivariate Gaussians are entirely characterised by their first two moments. In machine learning applications all that will ever be observed are evaluations of $f$ at finitely many points. Therefore we would like to study sample path regularity of $f$ up to modification. By this we mean that $f\sim\mathcal{GP}(\mu,k)$ will be said to be sample $\mathcal{F}(O)$, where $\mathcal{F}(O)$ is some space of functions $O\to\mathbb{R}$, if there exists a construction of the GP $f$, say $f\colon O\times\Omega\to\mathbb{R}$, such that $f(\cdot,\omega)\in\mathcal{F}(O)$ for all $\omega\in\Omega$.
Finally, note that $f\sim\mathcal{GP}(\mu,k)$ being sample $\mathcal{F}(O)$ is equivalent to $\tilde{f}+\mu$ being sample $\mathcal{F}(O)$, where $\tilde{f}\sim\mathcal{GP}(0,k)$. For this it suffices that $\mu\in\mathcal{F}(O)$ and $\tilde{f}$ is sample $\mathcal{F}(O)$. Therefore, in what follows, we will consider only centered GPs.
Thus we are studying sample path regularity of GPs from the covariance kernel, and our goal is to link the regularity of the kernel to the regularity of the GP sample paths.
2.1 Setup
We study GPs defined on open subsets of Euclidean spaces: $O\subset\mathbb{R}^{d}$ is open, for some $d\in\mathbb{N}$, and $f\sim\mathcal{GP}(0,k)$ is a centered GP on $O$. An important special case is when $k$ is stationary, i.e. $O=\mathbb{R}^{d}$ and there is an even function $k_{\delta}\colon\mathbb{R}^{d}\to\mathbb{R}$ such that
$k(\bm{x},\bm{y})=k_{\delta}(\bm{x}-\bm{y})$ (1)
for all $\bm{x},\bm{y}\in\mathbb{R}^{d}$. Another important case is when $k$ is isotropic, i.e. $O=\mathbb{R}^{d}$ and there is an even function $k_{r}\colon\mathbb{R}\to\mathbb{R}$ such that
$k(\bm{x},\bm{y})=k_{\delta}(\bm{x}-\bm{y})=k_{r}(\|\bm{x}-\bm{y}\|)$ (2)
[…]
3 Hölder regularity
The main type of sample path regularity we would like to investigate is continuity and the order of continuous differentiability. These are encapsulated in the framework of Hölder regularity.
We choose $O$ to be an open set in $\mathbb{R}^{d}$ in order to capture differentiability of the sample paths. However this implies that Hölder spaces on $O$ are inconvenient for our purposes, as they do not solely characterise local regularity properties of functions on $O$. Indeed, the Hölder spaces not only constrain the local behaviour of functions, but also their behaviour at infinity, or near the boundary of $O$. To retain solely local constraints we employ local Hölder spaces.
Definition 6 (Local Hölder and almost-Hölder spaces)
Let $n\in\mathbb{N}_{0}$ and $\gamma\in[0,1]$.
  1. (1)
    $C^{n,\gamma}_{loc}(O)$ is the space of functions $f$ on $O$ for which $\partial^{\bm{\alpha}}f$ exists for all multi-indices $\bm{\alpha}=(\alpha_{1},\dots,\alpha_{d})\in\mathbb{N}_{0}^{d}$ with $|\bm{\alpha}|:=\alpha_{1}+\dots+\alpha_{d}\leq n$, and such that the highest order partial derivatives satisfy a Hölder condition of the form: for all compact subsets $K\subset O$ there is a constant $C_{K}>0$ such that
    $|\partial^{\bm{\alpha}}f(\bm{x})-\partial^{\bm{\alpha}}f(\bm{y})|\leq C_{K}\|\bm{x}-\bm{y}\|^{\gamma}$
    for all $\bm{x},\bm{y}\in K$ and $|\bm{\alpha}|=n$.
  2. (2)
    $C^{(n+\gamma)^{-}}_{loc}(O):=\bigcap_{n^{\prime}+\gamma^{\prime}<n+\gamma}C^{n^{\prime},\gamma^{\prime}}_{loc}(O)$, assuming $n+\gamma>0$.
Remark 7
For $n,n^{\prime}\in\mathbb{N}_{0}$ and $\gamma,\gamma^{\prime}\in[0,1]$, $n^{\prime}+\gamma^{\prime}<n+\gamma$ implies $C^{n^{\prime},\gamma^{\prime}}_{loc}(O)\supsetneq C^{(n+\gamma)^{-}}_{loc}(O)\supsetneq C^{n,\gamma}_{loc}(O)$. Moreover $C^{n}(O)=C^{n,0}_{loc}(O)$.
As will become clear, the local almost-Hölder spaces in Definition 6(2) are natural for characterising GP sample path regularity.
To describe the regularity of the covariance kernel $k$, we will require derivatives of the form $\partial^{\bm{\alpha},\bm{\alpha}}k$. Here $\partial^{\bm{\alpha},\bm{\alpha}}$ stands for the successive application of $\partial^{\bm{\alpha}}$ with respect to the first variable and to the second variable. In fact the order in which they are applied will not matter, by continuity of these partial derivatives.
To obtain a certain almost-Hölder sample path regularity for $f$, we require the existence of the corresponding derivatives on the kernel applied to both variables. In addition we require an almost-Hölder continuity assumption on the highest order derivatives at the diagonal $\operatorname{diag}(O\times O):=\{(\bm{x},\bm{x}):\bm{x}\in O\}$, of an exponent equal to twice the Hölder exponent of the sample paths. To express this, recall the “big O” notation: the functions $f,g\colon\mathbb{R}^{d}\setminus\{\bm{0}\}\to\mathbb{R}$ satisfy $f(\bm{h})=\mathcal{O}(g(\bm{h}))$ as $\bm{h}\to\bm{0}$ if and only if there is $C>0$ such that $\limsup_{\bm{h}\to\bm{0}}\left|f(\bm{h})/g(\bm{h})\right|\leq C$. Moreover, for a family of functions $(f_{\bm{x}})_{\bm{x}\in\mathbb{R}^{l}}$, we say $f_{\bm{x}}(\bm{h})=\mathcal{O}(g(\bm{h}))$ as $\bm{h}\to\bm{0}$ locally uniformly in $\bm{x}\in\mathbb{R}^{l}$ if, for every compact subset $K\subset\mathbb{R}^{l}$, $C>0$ may be chosen independently of $\bm{x}\in K$.
We are now in a position to state the main result of this paper. {restatable}[Sample path Hölder regularity]theoremthmholder Let $n\in\mathbb{N}_{0}$ and $\gamma\in(0,1]$. The process $f\sim\mathcal{GP}(0,k)$ is sample $C^{(n+\gamma)^{-}}_{loc}(O)$ if and only if,
  1. (1)
    for general $k$,
    • $\partial^{\bm{\alpha},\bm{\alpha}}k$ exists and is continuous for all $|\bm{\alpha}|\leq n$, and
    • $|\partial^{\bm{\alpha},\bm{\alpha}}k(\bm{x}+\bm{h},\bm{x}+\bm{h})-2\partial^{\bm{\alpha},\bm{\alpha}}k(\bm{x}+\bm{h},\bm{x})+\partial^{\bm{\alpha},\bm{\alpha}}k(\bm{x},\bm{x})|=\mathcal{O}(\|\bm{h}\|^{2\epsilon})$ as $\bm{h}\to\bm{0}$
      locally uniformly in $\bm{x}\in O$, for all $\epsilon\in(0,\gamma)$ and $|\bm{\alpha}|=n$.
  2. (2)
    for stationary $k(\bm{x},\bm{y})=k_{\delta}(\bm{x}-\bm{y})$,
    • $\partial^{2\bm{\alpha}}k_{\delta}$ exists and is continuous for all $|\bm{\alpha}|\leq n$, and
    • $|\partial^{2\bm{\alpha}}k_{\delta}(\bm{h})-\partial^{2\bm{\alpha}}k_{\delta}(\bm{0})|=\mathcal{O}(\|\bm{h}\|^{2\epsilon})$ as $\bm{h}\to\bm{0}$ for all $\epsilon\in(0,\gamma)$ and $|\bm{\alpha}|=n$.
  3. (3)
    for isotropic $k(\bm{x},\bm{y})=k_{r}(\|\bm{x}-\bm{y}\|)$,
    • $k_{r}\in C^{2n}(\mathbb{R})$, and
    • $|k_{r}^{(2n)}(h)-k_{r}^{(2n)}(0)|=\mathcal{O}(|h|^{2\epsilon})$ as $h\to 0$ for all $\epsilon\in(0,\gamma)$.
In each case, differentiating sample-wise we have $\partial^{\bm{\alpha}}f\sim\mathcal{GP}(0,\partial^{\bm{\alpha},\bm{\alpha}}k)$ for all $|\bm{\alpha}|\leq n$. The proof can be found in Appendix A.
Remark 8
The positive definiteness of $k$ allows us to extrapolate regularity: $k$, $k_{\delta}$ or $k_{r}$ have no more derivatives around the diagonal, $\bm{0}$ or 00 respectively as they do on the rest of the domain. Therefore, when applying Theorem 7, it suffices to check the existence of the respective derivatives around the diagonal, $\bm{0}$ or 00 respectively.
Theorem 7 combines results about sample Hölder continuity of GPs (Azmoodeh et al., 2014) to results about sample differentiability of GPs (Potthoff, 2010), applied inductively on the partial derivatives. It gives us necessary and sufficient conditions for $f$ to be sample almost-Hölder continuous to a certain degree (i.e. $C^{(n+\gamma)^{-}}_{loc}(O)$ for some $n$, $\gamma$). Let us note however that it does not quite give us necessary conditions for $f$ to be sample $C^{n}(O)$, as the latter is not an almost-Hölder space. To achieve this, we can adapt the proof of Theorem 7 combining necessary and sufficient conditions for sample continuity of GPs with the results about sample differentiability from Potthoff (2010) (see Scheuerer (2010a, Theorem 5.3.16) for a result of this flavour). Characterising necessary and sufficient conditions on the kernel for the sample continuity of the GP is more involved than almost-Hölder continuity, and therefore we do not expand our results to this setting (see Fernique (1975) for the stationary case). Moreover we demonstrate in the examples below that Theorem 7 is all we need.
4 Examples
In this section we demonstrate how Theorem 7 can be applied in practice. We investigate the sample path regularity of the most widely used non-smooth GP families, recovering known results as well as proving novel ones.
By the reverse implications in Theorem 7, all the sample path regularity results in this section are sharp in the following sense: if $f$ is said to be sample $C^{(n+\gamma)^{-}}_{loc}(O)$ then $f$ is not sample $C^{(n^{\prime}+\gamma^{\prime})^{-}}_{loc}(O)$ for any $n^{\prime}+\gamma^{\prime}>n+\gamma$.
4.1 Wiener Gaussian process
The Wiener process is a centered GP on $O=\mathbb{R}_{>0}$ with covariance kernel
$k(x,y)=\min(x,y)$
for $x,y\in\mathbb{R}_{>0}$. $\min(\cdot,\cdot)$ is Lipschitz but non-differentiable in both its arguments. So by Theorem 7(1) the Wiener process is sample $C^{1/2^{-}}_{loc}(\mathbb{R}_{>0})$.
4.2 Matérn Gaussian processes
The Matérn kernels are isotropic kernels on $\mathbb{R}^{d}$ given by
$k(\bm{x},\bm{y})=k_{r}(\|\bm{x}-\bm{y}\|)=\frac{2^{1-\nu}}{\Gamma(\nu)}\left(\sqrt{2\nu}\|\bm{x}-\bm{y}\|\right)^{\nu}K_{\nu}\left(\sqrt{2\nu}\|\bm{x}-\bm{y}\|\right)$ (3)
for $\bm{x},\bm{y}\in\mathbb{R}^{d}$, where $K_{\nu}$ is the modified Bessel function of the second kind, and $\nu>0$ is the smoothness parameter. The cases where $\nu=n+1/2$ for some $n\in\mathbb{N}_{0}$ are particularly interesting because the expression in Equation (3) simplifies to a product of a polynomial and an exponential of the radial distance (Rasmussen and Williams, 2005, Equation 4.16).
Proposition 9
A centered Matérn GP with smoothness parameter $\nu$ is sample $C^{\nu^{-}}_{loc}(O)$. In particular, if $\nu=n+1/2$ for some $n\in\mathbb{N}_{0}$, the GP is sample $C^{(n+1/2)^{-}}_{loc}(\mathbb{R}^{d})$.
Proof  For $\rho>0$ we can write
$K_{\nu}(\rho)=\frac{\pi}{2}\frac{I_{-\nu}(\rho)-I_{\nu}(\rho)}{\sin(\nu\pi)}$
where
$I_{\pm\nu}(\rho)=\left(\frac{\rho}{2}\right)^{\pm\nu}\sum_{n=0}^{\infty}\frac{1}{n!\,\Gamma(n\pm\nu+1)}\left(\frac{\rho}{2}\right)^{2n}$
are modified Bessel functions of the first kind (Abramowitz and Stegun, 1965, Equations 9.6.2 & 9.6.10). So we can write the Matérn kernel (3) as
$k_{r}(x)=C_{\nu}\bigg{(}\underbrace{\sum_{n=0}^{\infty}\frac{1}{n!\,\Gamma(n-\nu+1)}\left(\frac{x}{2}\right)^{2n}}_{(*)}-\frac{|x|^{2\nu}}{2^{\nu}}\underbrace{\sum_{n=0}^{\infty}\frac{1}{n!\,\Gamma(n+\nu+1)}\left(\frac{x}{2}\right)^{2n}}_{(**)}\bigg{)}$
for $x\neq 0$, where $C_{\nu}>0$ is some constant. The power series (*)(*)( * ) and (**)(**)( * * ) are smooth in $x$, and $|x|^{2\nu}$ is $2\lfloor\nu\rfloor$ times continuously differentiable around 00 with $(2\nu-2\lfloor\nu\rfloor)$-Hölder continuous $2\lfloor\nu\rfloor$th derivative at 0. So the result follows by Theorem 7(3).
Remark 10
Proposition 9 also covers the (multivariate) Ornstein-Uhlenbeck process, the centered GP with isotropic covariance kernel given by
$k(\bm{x},\bm{y})=k(\|\bm{x}-\bm{y}\|)=\exp(-\|\bm{x}-\bm{y}\|)$
which corresponds to the Matérn GP with $\nu=1/2$. So the (multivariate) Ornstein-Uhlenbeck process is sample $C^{1/2^{-}}_{loc}(\mathbb{R}^{d})$.
4.3 Wendland Gaussian processes
The Wendland kernels are isotropic kernels which are compactly supported piecewise polynomials in the radial distance. On $\mathbb{R}^{d}$, they are defined as follows (Wendland, 2004, Definition 9.11):
$k(\bm{x},\bm{y})=k_{r}(\|\bm{x}-\bm{y}\|)=\mathcal{I}^{n}\phi_{\lfloor d/2\rfloor+n+1}(\|\bm{x}-\bm{y}\|)$
[…]
Proposition 11
A centered Wendland GP with degree parameter $n$ is sample $C^{(n+1/2)^{-}}_{loc}(O)$.
Proof  It is shown in (Wendland, 2004, Theorem 9.12) that the Wendland kernels may be written as
$k_{r}(x)=\sum_{j=0}^{\lfloor d/2\rfloor+3n+1}d^{(\lfloor d/2\rfloor+n+1)}_{j,n}|x|^{j}$
for $x\in\mathbb{R}$, where $d^{(\lfloor d/2\rfloor+n+1)}_{j,n}\in\mathbb{R}$ are coefficients. Furthermore the odd degree coefficients satisfy $d^{(\lfloor d/2\rfloor+n+1)}_{2j+1,n}=0$ if and only if $10\leq j\leq 2n-$ 1. Now for all $j\in\mathbb{N}_{0}$, $|x|^{2j}$ is smooth and $|x|^{2j+1}$ is $2n$ times continuously differentiable with Lipschitz $2n$th derivative. So the result follows by Theorem 7(3).
4.4 Tensor product Gaussian processes
GPs with tensor product covariance kernels are interesting as they allow spherically asymmetric sample path regularity. To formalise what is meant by this, let $O_{i}\subset\mathbb{R}^{d_{i}}$ be open sets for $1\leq i\leq l$ and $O_{1}\times\dots\times O_{l}=:O\subset\mathbb{R}^{d}=\mathbb{R}^{d_{1}}\times\dots\times\mathbb{R}^{d_{l}}$.
Definition 12
Let $n_{1},\dots,n_{l}\in\mathbb{N}_{0}$ and $\gamma_{1},\dots,\gamma_{l}\in[0,1]$.
  1. (1)
    $C^{(n_{1},\gamma_{1})\otimes\dotsb\otimes(n_{l},\gamma_{l})}_{loc}(O_{1}\times\dotsb\times O_{l})$ is the space of functions $f$ on $O$ for which $\partial^{\bm{\alpha}}f$ exists and is continuousauthor=Marvin,color=yellowauthor=Marvin,color=yellowtodo: author=Marvin,color=yellowdo we need this here? for all multi-indices $\bm{\alpha}\in\mathbb{N}_{0}^{d}$ with $|\bm{\alpha}_{i}|\leq n_{i}$, where $\bm{\alpha}_{i}\in\mathbb{N}_{0}^{d_{i}}$ is the projection of $\bm{\alpha}$ onto $\mathbb{N}_{0}^{d_{i}}$, and such that the highest order partial derivatives satisfy a Hölder condition of the form: for all compact subsets $K\subset O$ there is a constant $C_{K}>0$ such that
    $|\partial^{\bm{\alpha}}f(\bm{x})-\partial^{\bm{\alpha}}f(\bm{y})|\leq C_{K}\|\bm{x}_{i}-\bm{y}_{i}\|^{\gamma_{i}}$
    for all $1\leq i\leq d$, $\bm{x},\bm{y}\in K$ with $\bm{x}_{j}=\bm{y}_{j}$ for $j\neq i$, and $|\bm{\alpha}_{i}|=n_{i}$.author=Marvin,color=yellowauthor=Marvin,color=yellowtodo: author=Marvin,color=yellowWe might be able to get away with $|\alpha_{j}|=n_{j}$ for all $1\leq j\leq d$.
  2. (2)
    $C^{(n_{1}+\gamma_{1})^{-}\otimes\dotsb\otimes(n_{l}+\gamma_{l})^{-}}_{loc}(O_{1}\times\dotsb\times O_{l})=\bigcap_{n_{i}^{\prime}+\gamma_{i}^{\prime}<n_{i}+\gamma_{i}}C^{(n_{1}^{\prime},\gamma_{1}^{\prime})\otimes\dotsb\otimes(n_{l}^{\prime},\gamma_{l}^{\prime})}_{loc}(O_{1}\times\dotsb\times O_{l})$, assuming $n_{i}+\gamma_{i}>0$ for all $1\leq i\leq d$.
If $k_{i}\colon O_{i}\times O_{i}\to\mathbb{R}$ are symmetric positive definite kernels for $1\leq i\leq l$, then the tensor product kernel $k\colon O\times O\to\mathbb{R}$, given by
$k(\bm{x},\bm{y}):=(k_{1}\otimes\dotsb\otimes k_{l})(\bm{x},\bm{y}):=\prod_{i=1}^{l}k_{i}(\bm{x}_{i},\bm{y}_{i})$
for $\bm{x},\bm{y}\in\mathbb{R}^{d}$. The following proposition generalises Theorem 7.
Proposition 13
For $n_{1},\dots,n_{l}\in\mathbb{N}_{0}$ and $\gamma_{1},\dots,\gamma_{l}\in(0,1]$, the process $f\sim\mathcal{GP}(0,k)$ is sample $C^{(n_{1}+\gamma_{1})^{-}\otimes\dots\otimes(n_{l}+\gamma_{l})^{-}}_{loc}(O_{1}\times\dotsb\times O_{l})$ if and only if the $k_{i}$ satisfy the condition in Theorem 7(1) for all $1\leq i\leq l$.
The proof of Proposition 13 follows the same steps as the proof of Theorem 7, see Appendix A.
Remark 14
We could generalise Theorem 7 to allow for even more flexibility in the asymmetrical derivatives in Theorem 7. This can be done by defining Hölder spaces extending the spaces $C^{A}(O)$, where $A\subset\mathbb{N}_{0}^{d}$ is a downward closed set of multi-indices which specifies which partial derivatives exist (Pförtner et al., 2023, Definition B.9).
5 Sobolev regularity
In this section we investigate how the regularity of the kernel affects the weak differentiability of the sample paths. Specifically, we consider the important $L^{2}$-Sobolev regularity. The more general $L^{p}$-Sobolev regularity is studied in Henderson (2024). To characterise only the local regularity of the sample paths, as for Hölder spaces, we define the local pre-Sobolev spaces.
Definition 15 (Local pre-Sobolev spaces)
Let $n\in\mathbb{N}_{0}$. $\mathcal{H}^{n}_{loc}(O)$ is the space of functions [^1] A (local) Sobolev space is an $L^{2}$ quotient of a (local) pre-Sobolev space. Since we are interested in function spaces, pre-Sobolev spaces are the right framework for us. $f$ on $O$ for which for every compact $K\subset O$ the $L^{2}$ weak derivative $\partial^{\bm{\alpha}}_{w}f$ exists on $K$ for all multi-index $\bm{\alpha}\in\mathbb{N}_{0}^{d}$ with $|\bm{\alpha}|\leq n$, i.e. for all such $\bm{\alpha}$ there is a function $\partial^{\bm{\alpha}}_{w}f\in L^{2}(K)$ such that
$\int_{K}\partial^{\bm{\alpha}}_{w}f(\bm{x})\varphi(\bm{x})\,d\bm{x}=(-1)^{|\bm{\alpha}|}\int_{K}f(\bm{x})\partial^{\bm{\alpha}}\varphi(\bm{x})\,d\bm{x}$
for all $\varphi\in C^{\infty}(K)$.
Theorem 16 (Sample path Sobolev regularity)
Let $n\in\mathbb{N}$. The process $f\sim\mathcal{GP}(0,k)$ is sample $\mathcal{H}^{n}_{loc}(O)$ if,
  1. (1)
    for general $k$, $\partial^{\bm{\alpha},\bm{\alpha}}k$ exists and is continuous for all $|\bm{\alpha}|\leq n$.
  2. (2)
    for stationary $k(\bm{x},\bm{y})=k_{\delta}(\bm{x}-\bm{y})$, $\partial^{2\bm{\alpha}}k_{\delta}$ exists at $\bm{0}$ for all $|\bm{\alpha}|\leq n$.
  3. (3)
    for isotropic $k(\bm{x},\bm{y})=k_{r}(\|\bm{x}-\bm{y}\|)$, $k_{r}^{(j)}$ exists at 00 for all $j\leq 2n$.
If we assume that $k$ is continuous, or more generally that $f$ has a measurable modification, then the converse holds in the scenarios (2) and (3).
Proof  (1) follows from the general theorem for second order measurable random fields (Scheuerer, 2010b, Theorem 1). (2) and (3) follow from its corollary (Scheuerer, 2010b, Corollary 1). The converse in (2) follows from (Scheuerer, 2010b, Proposition 1). Finally, if (3) holds then, by the propagation of regularity from Gneiting (1999), we have $k_{r}\in C^{2n}(\mathbb{R})$. So by Lemma A.1 (3) $\Rightarrow$ (2) (without the Hölder condition) we deduce (2), and hence we obtain the converse statement in this case too.
As observed in Scheuerer (2010b), Sobolev differentiability is a natural regularity notion for sample paths of general random fields, as it can be deduced from continuous mean square differentiability. However, in the case of the GPs, Hölder control on the highest order partial derivatives of the kernel at the diagonal is sufficient to deduce strong differentiability of the sample paths (Theorem 7), and not merely weak differentiability. In other words, Theorem 16 can be seen as a negative result, and we expect it to have less practical value for GPs as its general version (Scheuerer, 2010b, Theorem 1) for second order measurable random fields. Note that deducing sample path Hölder regularity from Theorem 16 followed by Sobolev embedding theorems introduces a superfluous dependence on dimension, and yields weaker results than the dimension independent Theorem 7.
Acknowledgments and Disclosure of Funding
NDC, MP and PH gratefully acknowledge financial support by the European Research Council through ERC StG Action 757275 / PANAMA; the DFG Cluster of Excellence “Machine Learning - New Perspectives for Science”, EXC 2064/1, project number 390727645; the German Federal Ministry of Education and Research (BMBF) through the Tübingen AI Center (FKZ: 01IS18039A); and funds from the Ministry of Science, Research and Arts of the State of Baden-Württemberg. NDC acknowledges the support of the Fonds National de la Recherche, Luxembourg. The authors thank the International Max Planck Research School for Intelligent Systems (IMPRS-IS) for supporting NDC and MP.
Appendix A Proof of Theorem 7
We start by showing the equivalences of the various kernel conditions in Theorem 7, namely (1) $\Leftrightarrow$ (2) in the stationary case, and (1) $\Leftrightarrow$ (2) $\Leftrightarrow$ (3) in the isotropic case.
Lemma A.1
Let $n\in\mathbb{N}_{0}$ and $\epsilon\in(0,1]$. Then the condition
  1. (1)
    • $\partial^{\bm{\alpha},\bm{\alpha}}k$ exists and is continuous for all $|\bm{\alpha}|\leq n$, and
    • $|\partial^{\bm{\alpha},\bm{\alpha}}k(\bm{x}+\bm{h},\bm{x}+\bm{h})-2\partial^{\bm{\alpha},\bm{\alpha}}k(\bm{x}+\bm{h},\bm{x})+\partial^{\bm{\alpha},\bm{\alpha}}k(\bm{x},\bm{x})|=\mathcal{O}(\|\bm{h}\|^{2\epsilon})$ as $\bm{h}\to\bm{0}$
      locally uniformly in $\bm{x}\in O$ for all $|\bm{\alpha}|=n$.
is equivalent to the following:
  1. (2)
    for stationary $k(\bm{x},\bm{y})=k_{\delta}(\bm{x},\bm{y})$,
    • $\partial^{2\bm{\alpha}}k_{\delta}$ exists and is continuous for all $|\bm{\alpha}|\leq n$, and
    • $|\partial^{2\bm{\alpha}}k_{\delta}(\bm{h})-\partial^{2\bm{\alpha}}k_{\delta}(\bm{0})|=\mathcal{O}(\|\bm{h}\|^{2\epsilon})$ as $\bm{h}\to\bm{0}$ for all $|\bm{\alpha}|=n$.
  2. (3)
    for isotropic $k(\bm{x},\bm{y})=k_{r}(\|\bm{x}-\bm{y}\|)$,
    • $k_{r}\in C^{2n}(\mathbb{R})$, and
    • $|k_{r}^{(2n)}(h)-k_{r}^{(2n)}(0)|=\mathcal{O}(|h|^{2\epsilon})$ as $h\to 0$.
Proof of Lemma A.1 (1) normal-$\Leftrightarrow$ (2) Let $k$ be stationary, i.e. $k(\bm{x},\bm{y})=k_{\delta}(\bm{x}-\bm{y})$.
(2) normal-$\Rightarrow$ (1): For $|\bm{\alpha}|\leq n$,
$\partial^{\bm{\alpha},\bm{\alpha}}k(\bm{x},\bm{y})=(-1)^{|\bm{\alpha}|}\partial^{2\bm{\alpha}}k_{\delta}(\bm{x}-\bm{y})$
for all $\bm{x},\bm{y}\in\mathbb{R}^{d}$, by the chain rule. So the existence and the continuity of $\partial^{2\bm{\alpha}}k_{\delta}$ imply the existence and the continuity of $\partial^{\bm{\alpha},\bm{\alpha}}k$. Moreover when $|\bm{\alpha}|=n$,
$|\partial^{\bm{\alpha},\bm{\alpha}}k(\bm{x}+\bm{h},\bm{x}+\bm{h})-2\partial^{\bm{\alpha},\bm{\alpha}}k(\bm{x}+\bm{h},\bm{x})+\partial^{\bm{\alpha},\bm{\alpha}}k(\bm{x},\bm{x})|$ (4)
$=2|\partial^{2\bm{\alpha}}k_{\delta}(\bm{0})-\partial^{2\bm{\alpha}}k_{\delta}(\bm{h})|,$
for all $\bm{x},\bm{h}\in\mathbb{R}^{d}$, so the Hölder conditions on the highest order partial derivatives of $k$ at the diagonal and of $k_{\delta}$ at $\bm{0}$ correspond.
(1) normal-$\Rightarrow$ (2): For $|\bm{\alpha}|\leq n$ and $\bm{x}\in\mathbb{R}^{d}$, $\partial^{2\bm{\alpha}}k_{\delta}(\bm{x})$ exists and is equal to $(-1)^{|\bm{\alpha}|}\partial^{\bm{\alpha},\bm{\alpha}}k(\bm{x},\bm{0})$. Moreover the Hölder conditions on the derivatives of $k$ and $k_{\delta}$ correspond by Equation (4).
To prove the second equivalence, we will need the following result:
Lemma A.2
Let $m\in\mathbb{N}_{0}$, $g\in C^{m}(\mathbb{R}\setminus\{0\})$, and $\nu\in\mathbb{R}$ such that $g^{(j)}(h)=\mathcal{O}(|h|^{\nu-j})$ as $h\to 0$, for all $0\leq j\leq m$. Define $f:=g\circ\|\cdot\|\colon\mathbb{R}^{d}\setminus\{\bm{0}\}\to\mathbb{R}$. Then $f\in C^{m}(\mathbb{R}^{d}\setminus\{\bm{0}\})$ and $\partial^{\bm{\alpha}}f(\bm{h})=\mathcal{O}(\|\bm{h}\|^{\nu-j})$ as $\bm{h}\to\bm{0}$, for $0\leq j\leq m$ and $|\bm{\alpha}|=j$.
Proof  By induction on $m$. $m=0$ is clear; the growth/decay of $f$ at $\bm{0}$ is the same as that of $g$ at 00. For $m>0$, $\bm{x}\in\mathbb{R}^{d}\setminus\{\bm{0}\}$, and $1\leq i\leq d$, note that
$\partial^{\bm{e}_{i}}f(\bm{x})=g^{\prime}(\|\bm{x}\|)\frac{x_{i}}{\|\bm{x}\|}=\tilde{g}(\|\bm{x}\|)x_{i}$ (5)
where $\bm{e}_{i}$ is the $i$th unit vector and $\tilde{g}(x):=\frac{g^{\prime}(x)}{x}$ for $x\in\mathbb{R}\setminus\{0\}$. Clearly $\tilde{g}\in C^{m-1}(\mathbb{R}\setminus\{0\})$ and
$\tilde{g}^{(j)}(h)=\sum_{l=0}^{j}\binom{j}{l}g^{(l+1)}(h)(-1)^{j-l}(j-l)!h^{-1-j+l}=\mathcal{O}(|h|^{(\nu-2)-j})$
as $h\to 0$, for all $10\leq j\leq m-$ 1. Let $\tilde{f}=\tilde{g}\circ\|\cdot\|$. By the induction hypothesis, $\partial^{\bm{\alpha}}\tilde{f}(\bm{h})=\mathcal{O}(\|\bm{h}\|^{(\nu-2)-j})$ as $\bm{h}\to\bm{0}$, for $|\bm{\alpha}|=j$ and $10\leq j\leq m-$ 1. Thus, by (5),
$\partial^{\bm{\alpha}}\partial^{\bm{e}_{i}}f(\bm{h})=\left\{\begin{aligned} &h_{i}\partial^{\bm{\alpha}}\tilde{f}(\bm{h})&\ &\text{if }\alpha_{i}=0\\ &h_{i}\partial^{\bm{\alpha}}\tilde{f}(\bm{h})+\partial^{\bm{\alpha}-\bm{e}_{i}}\tilde{f}(\bm{h})&&\text{if }\alpha_{i}>0\end{aligned}\right\}=\mathcal{O}(\|\bm{h}\|^{\nu-(j+1)})$
as $\bm{h}\to\bm{0}$. Consequently, $\partial^{\bm{\alpha}}f(\bm{h})=\mathcal{O}(\|\bm{h}\|^{\nu-j})$ as $\bm{h}\to\bm{0}$ for $|\bm{\alpha}|=j$ and $0\leq j\leq m$. This concludes the induction step.
Proof of Lemma A.1 (1) normal-$\Leftrightarrow$ (3) Let $k$ be isotropic, i.e. $k(\bm{x},\bm{y})=k_{r}(\|\bm{x}-\bm{y}\|)$. By Lemma A.1 (1) $\Leftrightarrow$ (2), it suffices to show that the condition on $k_{\delta}$ in (2) is equivalent to the condition on $k_{r}$ in (3).
(2) normal-$\Rightarrow$ (3): This follows from the fact that $x\mapsto k_{\delta}(x\bm{e}_{1})$ is exactly $k_{r}$, and analogously $x\mapsto\partial^{2j\bm{e}_{1}}k_{\delta}(x\bm{e}_{1})$ is equal to $k_{r}^{(2j)}$ for $0\leq j\leq n$.
(3) normal-$\Rightarrow$ (2): In our context $k_{r}\colon\mathbb{R}\to\mathbb{R}$ is even. So the odd derivatives of $k_{r}$ which exist at 0 vanish there. Let
$g(x):=k_{r}(x)-\sum_{j=0}^{n}\frac{x^{2j}}{j!}k_{r}^{(2j)}(0),$
for all $x\in\mathbb{R}$. The Lagrange form of the remainder in an order $12n-j-$ 1 Taylor expansion of $g^{(j)}$ then reveals that
$g^{(j)}(h)=\frac{h^{2n-j}}{(2n-j)!}(k_{r}^{(2n)}(\xi_{j}(h))-k_{r}^{(2n)}(0))=\mathcal{O}(|h|^{2n-j+2\epsilon})$
[…]
$k_{\delta}(\bm{x})=f(\bm{x})+\sum_{j=0}^{n}\frac{\|\bm{x}\|^{2j}}{j!}k_{r}^{(2j)}(0)$
is in $C^{2n}(\mathbb{R})$. Moreover, for $\bm{h}\in\mathbb{R}^{d}$ and $|\bm{\alpha}|=n$ there is a constant $C_{\bm{\alpha}}\in\mathbb{R}$ such that $\partial^{2\bm{\alpha}}k_{\delta}(\bm{h})=\partial^{2\bm{\alpha}}f(\bm{h})+C_{\bm{\alpha}}$, implying that
$|\partial^{2\bm{\alpha}}k_{\delta}(\bm{h})-\partial^{2\bm{\alpha}}k_{\delta}(\bm{0})|=|\partial^{2\bm{\alpha}}f(\bm{h})-\underbrace{\partial^{2\bm{\alpha}}f(\bm{0})}_{=0}|=\mathcal{O}(\|\bm{h}\|^{2n+2\epsilon-2n})=\mathcal{O}(\|\bm{h}\|^{2\epsilon})$
as $\bm{h}\to\bm{0}$. Thus $k_{\delta}$ satisfies the Hölder condition in (2).
Note that Lemma A.1 is purely a result about stationary and isotropic functions; the proof does make use of the positive definiteness of $k$. In fact, the positive definiteness of $k$ gives us additional information about its regularity, see Remark 8.
[…]
Proof  We only prove the general case (1); cases (2) and (3) then follow by taking intersections over $\epsilon\in(0,\gamma)$ in Lemma A.1.
[…]
For $n=1$, the existence and continuity of $\partial^{\bm{e}_{i},\bm{e}_{i}}k$ around the diagonal, where $1\leq i\leq d$ and $\bm{e}_{i}$ is the $i$th unit vector, implies that $f$ is mean-square differentiable in direction $i$ (Potthoff, 2010, Lemma 2.8). Write $\partial^{\bm{e}_{i}}_{ms}f\colon O\times\Omega\to\mathbb{R}$ for a mean-square partial derivative of $f$ in direction $i$. We show $\partial^{\bm{e}_{i}}_{ms}f\sim\mathcal{GP}(0,\partial^{\bm{e}_{i},\bm{e}_{i}}k)$. We need to check that for any finite set of points $\{\bm{x}_{1},\dots,\bm{x}_{N}\}\subset O$,
$(\partial^{\bm{e}_{i}}_{ms}f(\bm{x}_{1}),\dots,\partial^{\bm{e}_{i}}_{ms}f(\bm{x}_{N}))\sim\mathcal{N}\left(\bm{0},\left(\partial^{\bm{e}_{i},\bm{e}_{i}}k(\bm{x}_{p},\bm{x}_{q})\right)_{p,q=1}^{N}\right).$ (6)
$(\partial^{\bm{e}_{i}}_{ms}f(\bm{x}_{1}),\dots,\partial^{\bm{e}_{i}}_{ms}f(\bm{x}_{N}))$ is the limit in $L^{2}(\Omega;\mathbb{R}^{N})$ as $h\to 0$ of
$\left(\frac{f(\bm{x}_{p}+h\bm{e}_{i})-f(\bm{x}_{p})}{h}\right)_{p=1}^{N}$ (7)
$\sim\mathcal{N}\bigg{(}\bm{0},\bigg{(}$ $\frac{k(\bm{x}_{p}+h\bm{e}_{i},\bm{x}_{q}+h\bm{e}_{i})-k(\bm{x}_{p}+h\bm{e}_{i},\bm{x}_{q})-k(\bm{x}_{p},\bm{x}_{q}+h\bm{e}_{i})+k(\bm{x}_{p},\bm{x}_{q})}{h^{2}}\bigg{)}_{p,q=1}^{N}\bigg{)}.$
Now convergence in $L^{2}$ implies convergence in distribution, and we see that the multivariate normal distribution in Equation (7) converges to the multivariate normal distribution in Equation (6) as $h\to 0$. So this shows $\partial^{\bm{e}_{i}}_{ms}f\sim\mathcal{GP}(0,\partial^{\bm{e}_{i},\bm{e}_{i}}k)$. Now applying the case $n=0$ to $\partial^{\bm{e}_{i},\bm{e}_{i}}k$ we deduce that $\partial^{\bm{e}_{i}}_{ms}f$ is sample [^2] The mean-square partial derivative $\partial^{\bm{e}_{i}}_{ms}f(\bm{x})$ at a point $\bm{x}\in O$ is only well-defined almost surely on $\Omega$. But we still talk of continuous sample paths for the process $\partial^{\bm{e}_{i}}_{ms}f$, since we mean this up to modification, as described in Section 2. $C^{\gamma^{-}}_{loc}(O)$. This implies by (Potthoff, 2010, Theorem 3.2) that $f$ is sample differentiable in direction $i$, with $\partial^{\bm{e}_{i}}f(\bm{x})=\partial^{\bm{e}_{i}}_{ms}f(\bm{x})$ almost surely, for all $\bm{x}\in O$. Thus $\partial^{\bm{e}_{i}}f\sim\mathcal{GP}(0,\partial^{\bm{e}_{i},\bm{e}_{i}}k)$ and $\partial^{\bm{e}_{i}}f$ is sample $C^{\gamma^{-}}_{loc}(O)$. $1\leq i\leq d$ was arbitrary, so $f$ is sample $C^{(1+\gamma)^{-}}_{loc}(O)$. For $n>1$, we apply the same argument inductively on the partial derivatives.
$\Longrightarrow$: For $n=0$, Azmoodeh et al. (2014, Theorem 1) gives the converse to Kolmogorov’s theorem for GPs.
Now suppose $n=1$. Pick $1\leq i\leq d$. For any finite set of points $\{\bm{x}_{1},\dots,\bm{x}_{N}\}\subset O$, $(\partial^{\bm{e}_{i}}f(\bm{x}_{1}),\dots,\partial^{\bm{e}_{i}}f(\bm{x}_{N}))$ is the almost sure limit of centered multivariate Gaussians with distribution
$\mathcal{N}\bigg{(}\bm{0},\bigg{(}\frac{k(\bm{x}_{p}+h\bm{e}_{i},\bm{x}_{q}+h\bm{e}_{i})-k(\bm{x}_{p}+h\bm{e}_{i},\bm{x}_{q})-k(\bm{x}_{p},\bm{x}_{q}+h\bm{e}_{i})+k(\bm{x}_{p},\bm{x}_{q})}{h^{2}}\bigg{)}_{p,q=1}^{N}\bigg{)}$
(see Equation (7)), so is itself a centered multivariate Gaussian distribution. Hence $\partial^{\bm{e}_{i}}f$ is a GP.
Let $\bm{x},\bm{x}+h\bm{e}_{i}\in O$. By the mean value theorem, for each $\omega\in\Omega$ there is $\xi_{\bm{x},\omega}(h)$ between $\bm{x}$ and $\bm{x}+h\bm{e}_{i}$ such that
$\frac{f(\bm{x}+h\bm{e}_{i},\omega)-f(\bm{x},\omega)}{h}=\partial^{\bm{e}_{i}}f(\xi_{\bm{x},\omega}(h),\omega).$
Thus
$\left|\frac{f(\bm{x}+h\bm{e}_{i},\omega)-f(\bm{x},\omega)}{h}-\partial^{\bm{e}_{i}}f(\bm{x},\omega)\right|^{2}=|\partial^{\bm{e}_{i}}f(\xi_{\bm{x},\omega}(h),\omega)-\partial^{\bm{e}_{i}}f(\bm{x},\omega)|^{2}\leq C_{\omega}^{2}|h|^{\gamma}$
since $\gamma\in(0,2\gamma)$, for some constant $C_{\omega}$ depending on $\omega$ but not on $h$, assuming $h$ is small enough.
We have
[…]
for $h$ small enough. (**)<(**)$\infty$( * * ) < ∞ by Azmoodeh et al. (2014, Theorem 1), which we can apply since we showed that $\partial^{\bm{e}_{i}}f$ is a GP. Also (*$(*)\to 0$ as $h\to 0$. Hence
$\frac{f(\bm{x}+h\bm{e}_{i})-f(\bm{x})}{h}\xrightarrow{L^{2}}\partial^{\bm{e}_{i}}f(\bm{x})$
as $h\to 0$, i.e. $f$ is mean-square differentiable in direction $i$. So
$\frac{k(\bm{x}+h\bm{e}_{i},\bm{y})-k(\bm{x},\bm{y})}{h}=\left\langle\frac{f(\bm{x}+h\bm{e}_{i})-f(\bm{x})}{h},f(\bm{y})\right\rangle_{L^{2}}\to\langle\partial^{\bm{e}_{i}}f(\bm{x}),f(\bm{y})\rangle_{L^{2}}$
as $h\to 0$, for $\bm{x},\bm{y}\in O$. Hence $\partial^{\bm{e}_{i},\bm{0}}k$ exists. Furthermore,
$\frac{\partial^{\bm{e}_{i},\bm{0}}k(\bm{x},\bm{y}+h\bm{e}_{i})-\partial^{\bm{e}_{i},\bm{0}}k(\bm{x},\bm{y})}{h}$ $=\left\langle\partial^{\bm{e}_{i}}f(\bm{x}),\frac{f(\bm{y}+h\bm{e}_{i})-f(\bm{y})}{h}\right\rangle_{L^{2}}$
$\to\langle\partial^{\bm{e}_{i}}f(\bm{x}),\partial^{\bm{e}_{i}}f(\bm{y})\rangle_{L^{2}}$
as $h\to 0$. Hence $\partial^{\bm{e}_{i},\bm{e}_{i}}k$ exists and $\partial^{\bm{e}_{i}}f\sim\mathcal{GP}(0,\partial^{\bm{e}_{i},\bm{e}_{i}}k)$. To show the Hölder condition in (1) we apply the converse to Kolmogorov’s theorem for GPs (Azmoodeh et al., 2014, Theorem 1) to $\partial^{\bm{e}_{i},\bm{e}_{i}}k$.
Finally, for $n>1$ we apply the same argument inductively on the partial derivatives.